OpenStudy (anonymous):

The polar coordinates of a certain point are r=5.05 cm, @(theta)= 250 degrees. (a) Find its Cartesian coordinates x and y. (d) Find the polar coordinates of the points with Cartesian coordinates (3x, −3y). Show work.

OpenStudy (anonymous):

a. Polar->Cartesian \[x= rcos(\theta);y= rsin(\theta)\] r=5.05 , theta= 250 degrees \[x= rcos(\theta)\]\[=5.05cos(250)\]\[=-1.727\] \[y= rsin(\theta)\]\[= 5.05sin(250)\]\[=-4.745\] b.Cartesian->Polar (3x, −3y) \[r=\sqrt(x^2+y^2);\theta=\arctan(y/x)\] \[r=\sqrt(x^2+y^2)\]\[=\sqrt((3x)^2+(-3y)^2)\]\[=\sqrt(9x^2+9y^2)\]\[=\sqrt(9(x^2+y^2))\]\[=3\sqrt(x^2+y^2)\] \[\theta=\arctan(-3y/3x)\]\[\theta=\arctan(-y/x)\] Where the x and y points supposed to be (3,-3) or (3x,-3y)?

OpenStudy (anonymous):

*Were

OpenStudy (anonymous):

The last.

OpenStudy (anonymous):

Okay well then the answer above are for the last. That's as far as you can go :)